OpenStudy (anonymous):

A trapezoid shaped lot has an area of 162 squared meter. its height is 9m. what is the measure of the longer base id the shorter base is 15 m?

OpenStudy (anonymous):

 A = b1 + b2/2 x h

OpenStudy (anonymous):

162 m^2 = 15m + b2 /2 x 9

OpenStudy (anonymous):

yes and solve for b2 now.

OpenStudy (anonymous):

162 m^2= 67.5m .. is it right?